# CLAP: Unsupervised 3D Representation Learning for Fusion 3D Perception via Curvature Sampling and Prototype Learning

### v2

@@ 3.2 Neural Field and Differentiable Rendering
Inspired by the success of UniPAD [58], CLAP applies a differentiable rendering decoder with neural field to conduct the masked-and-reconstruction pre-training. Different from [58], an additional surface signed distance field loss is applied to better optimize the scene geometry. Here we first introduce Neural Field, which is the basis for camera images and point clouds rendering, and then discuss the differentiable rendering process on depth and RGB values.
@@
$\mathbf{f}_{\text{p}}=f^{\text{tri}}(\mathbf{p},\hat{\mathbf{P}}_{t}),$ (5)
where $f^{\text{tri}}$ is an built-in module implemented in Pytorch [37]. Taking the concatenation of location $\mathbf{p}$ and queried feature $\mathbf{f}_{\text{p}}$ as inputs, we predict the signed distance value $s\in\mathbb{R}$ [5, 31] and color value $c\in\mathbb{R}^{3}$ [49] at $\mathbf{p}$ with $f^{\text{SDF}}$ and $f^{\text{RGB}}$.
@@
where $f^{\text{RGB}}$ and $f^{\text{SDF}}$ are parameterized by Multi-layer Perceptron.
Differentiable Rendering. Similar to [33, 49], we first sample $N_{\text{L}}$ or $N_{\text{C}}$ rays at the LiDAR or camera sensor origin $\mathbf{o}$, each of which is described by its normalized direction $\mathbf{d}$ and $\mathbf{o}$. Next, we sample $N_{\text{ray}}$ points following [49] along each ray. Here each point along the ray can be interpreted by $\mathbf{p}=\mathbf{o}+r\mathbf{d}$, where $r$ is the range from the sensor origin to the point $\mathbf{p}$. Thus the sampled point set can be annotated by $\{\mathbf{p}_{n}=\mathbf{o}+r_{n}\mathbf{d}\}_{n=1}^{N_{\text{ray}}}$ and we predict the estimated signed distance value $s_{n}$ and color value $c_{n}$ for them with $f^{\text{RGB}}$ and $f^{\text{SDF}}$. Following [49], we estimate the occupancy value $\alpha_{n}$ for each sampled point,
@@
Here $\Phi_{h}(x)=(1+e^{-hx})^{-1}$ stands for the sigmoid function paired with a learnable scalar $h$. After that, we predict the accumulated transmittance $t_{n}$ similar to [49]
@@
Based on $t_{n}$, we compute an unbiased and occlusion-aware weight $w_{n}=t_{n}\alpha_{n}$ [49] and integrate all samples along the ray to predict the range $\tilde{r}$ or color $\tilde{c}$ along this ray,
@@
For the observed LiDAR points, it is evident that signed distance value at those points are 0. The final loss function for differentiable rendering is a combination of L-1 loss on range and color predictions and the surface SDF loss.
@@
Figure 3: Inspiration of Curvature Sampling. It can be found that areas with low curvature tend to be less informative, like road plane, while those with high curvature provide more information (objects like cars).
@@ 3.3 Curvature Sampling
In order to make joint unsupervised representation learning feasible, we have to make $N_{\text{L}}\ll N_{\text{P}}$ and $N_{\text{C}}\ll H\cdot W\cdot N_{\text{cam}}$. Intuitively, uniform sampling with depth can be used, same as “Memory-friendly Ray Sampling” in [58]. But due to the relatively small sample number compared to the raw inputs ($\sim\frac{1}{100}$), this sampling method brings little improvement against separate pre-training, which is contradictory to our motivation. Hence, we need to sample more informative part of the scene for $\mathcal{L}_{\text{rend}}$. As shown in Figure 3, we are inspired by the observation that surface with higher curvature (surface of a vehicle) generally contains more information as compared to that with lower curvature (road plane). Therefore, we propose the Curvature Sampling for effective sampling. For each point $\mathbf{p}$ in the LiDAR point cloud $\mathbf{P}$, we first estimate the surface normal by deriving the signed distance function with respect to $\mathbf{p}$
@@
where $\mathbf{n}\in\mathbb{R}^{3}$ is the predicted normal. Then we normalized $\mathbf{n}$ to get the direction of the normal $\tilde{\mathbf{n}}=\frac{\mathbf{n}}{||\mathbf{n}||_{2}}$. Here $||\cdot||_{2}$ is the L-2 norm. Next we apply the differential operator on $\tilde{\mathbf{n}}$ with respect to $\mathbf{p}$ and get $\mathbf{c}\in\mathbb{R}^{3}$.
@@
Then we estimate the geodesic curvature [48] of each point $\mathbf{p}_{n}$ in LiDAR point cloud by computing the norm of $\mathbf{c}_{n}$ and use it as the sampling weights $\omega_{n}$, that is $\omega_{n}=||\mathbf{c}_{n}||_{2}$. With $\omega_{n}$, we sample $N_{\text{L}}$ points with a Multinomial Sampler implemented in PyTorch [37] for differentiable rendering. For pixels on image plane, we project the LiDAR point cloud back to image planes with $\mathcal{T}$, assign $\omega_{n}$ of each point to the projected pixel and apply a gaussian blur kernel of size $K_{\text{gaus}}$ to densify the weights, with which we sample $N_{\text{C}}$ pixels for $\mathcal{L}_{\text{rend}}$. As the estimation of curvature here is noisy especially in the first few epochs, we set up a warm-up stage with $N_{\text{warmup}}$ epochs with uniform sampling and after $N_{\text{warmup}}$ epochs, Curvature Sampling is utilized.
@@ 3.4 Prototype Learning
Expectation-Maximization. In order to guide the learnable prototypes to represent parts of the environment, we propose an Expectation-Maximization [35] training scheme to optimize the prototypes. We first project the LiDAR embeddings $\hat{\mathbf{P}}$ and camera embeddings $\hat{\mathbf{I}}$ separately with two projection heads $f^{\text{proj}}_{\text{P}}$ and $f^{\text{proj}}_{\text{I}}$ to the same dimension as $\mathbf{K}$
$\dot{\mathbf{P}}=f^{\text{proj}}_{\text{P}}(\hat{\mathbf{P}})\ \ ,\ \ \dot{\mathbf{I}}=f^{\text{proj}}_{\text{I}}(\hat{\mathbf{I}}),$ (13)
@@
In the Expectation step, we compute the probability $\hat{\mathbf{S}}_{\text{P/I}}$ that each prototype is assigned to each embeddings by applying a softmax operation on $\mathbf{S}_{\text{P/I}}$. Then for Maximization step, we expect to maximize the probability of the assignment between one prototype to one specific part of the scene and this is the same as minimizing the entropy of the similarity matrix. Hence, the EM loss is computed as,
$\mathcal{L}_{\text{EM}}=-\frac{1}{N_{\text{3D}}N_{\text{K}}}\sum_{n=1}^{N_{\text{3D}}}\sum_{m=1}^{N_{\text{K}}}\{\hat{\mathbf{S}}_{\text{P}}^{n,m}log\hat{\mathbf{S}}_{\text{P}}^{n,m}+\hat{\mathbf{S}}_{\text{I}}^{n,m}log\hat{\mathbf{S}}_{\text{I}}^{n,m}\}$ (15)
Swapping Prototype Prediction. To further explore interaction between modalities, we detach $\mathbf{S}_{\text{P/I}}$ and apply sinkhorn algorithm [9] to make them approximate to double stochastic matrix in $N_{\text{sink}}$ iterations. We denote the updated matrix as codes $\mathbf{Q}_{\text{P/I}}\in\mathbb{R}^{N_{\text{3D}}\times N_{\text{K}}}$. The swapping prototype prediction loss is formulated with a temperature parameter $\tau$, inspired by [4],
@@
Gram Matrix Minimization. When training the randomly initialized prototypes, the network might learn a short cut with all prototypes being the same [4], which is called collapse. To avoid this, we estimate similarity between prototypes by the gram matrix $\mathbf{G}=\mathbf{K}\mathbf{K}^{\intercal}$ of prototypes $\mathbf{K}$, the dimension of which is $\mathbf{G}\in\mathbb{R}^{N_{\text{K}}\times N_{\text{K}}}$. Finally we minimize the average of the non-diagonal elements of $\mathbf{G}$ in order to avoid collapse
@@
Table 1: Results for fine-tuning on 5% of training set in NuScenes [3]. “Init.” means the way to initialize models. To avoid the case the performance difference is brought by accelerating convergence, we first increase the training iteration of training-from-scratch model (BEVFusion [30]) until we observe convergence, which leads to the “Rand.” row. Then we pre-train the backbones with ALSO [2], OCC-MAE [34], SLidR [38], PPKT [28], UniPAD [58] and CLAP. We initialize BEVFusion with the pre-trained weights and fine-tune the model with the same number iterations as “Rand.”. We provide mAP and NDS as an evaluation of the overall performance of different models and highlight the best mAP and NDS with bold font. We also indicate the performance improvement by green color. “C.V.”, “Mot.”, “Bic.”, “Ped.” and “T.C.” are abbreviations for Construction Vehicle, Motorcycle, Bicycle, Pedestrian and Traffic Cone. All the results here are in %.
4 Experiments
Unsupervised 3D representation learning for fusion perception aims to pre-train both LiDAR and camera encoders and initialize downstream models with the pre-trained weights to gain performance improvement in downstream tasks. In this section, we design extensive experiments on the popular autonomous driving dataset NuScenes [3] and Waymo [45] to demonstrate the effectiveness of CLAP. To begin with, we describe experiment setups in Section 4.1. Next, we show and analyze main results in Section 4.2. Finally, we provide ablation study and visualizations separately in Section 4.3 and 4.4.
4.1 Settings
Datasets. We use the popular autonomous driving dataset NuScenes [3] and Waymo [45] to evaluate the performance of CLAP. NuScenes [3] uses one roof LiDAR and six surrounding cameras to collect data. The LiDAR is a 32-beam Velodyne and collecting frequency is 20Hz. The frequency of camera capturing is 12Hz. [3] conducts the synchronization and provides paired data of LiDAR point cloud and camera images. The whole NuScenes dataset contains 1000 scenes collected in Boston and Singapore. Each scene lasts for around 20 seconds and there are a total of 5.5 hours data. Following convention practice from [3, 46], we divide the whole dataset into training set with 850 scenes and validation set with 150 scenes. Waymo [45] uses one top 64-beam LiDAR, 4 corner LiDARs and 5 surrounding cameras to collect point clouds and camera images. Following the same practice in [45, 46], the collected 1000 scenes are split into training set (798 scenes) and validation set (202 scenes). All pre-trainings are conducted on the training set without labels and we conduct downstream 3D object detection training in few-shot setting via uniform sampling of the training data (NuScenes [3] $5\%$ and Waymo [45] $1\%$).
Downstream 3D Object Detectors. For NuScenes [3], we select the SOTA 3D object detector called BEVFusion [30] and for Waymo [45], we use CenterPoint [59]. Both BEVFusion and CenterPoint are implemented in the popular code repository for autonomous driving perception called OpenPCDet [46]. For evaluation metrics, we use average precisions of various categories (APs), mean average precision (mAP) and NuScenes Detection Score (NDS) [3] for NuScenes [3] and mAP (mean accurate precisions) and mAPH (mean accurate precisions with headings) at different difficulty levels (Level-1 and 2) for Waymo [45]. We follow a similar setting in [14, 53] to gradually increase training iterations of the from-scratch model until convergence is observed. Here convergence means further increasing training iterations will not improve the performance. Then the number of training iterations is fixed for fine-tuning pre-trained models. This setting avoids the case that pre-training only accelerates convergence and makes sure that pre-training indeed improve the performance of downstream models, that is improving the sample efficiency of the downstream task.
Baseline Pre-training Method for Fusion Perception. We incorporate three kinds of pre-training baseline methods: 1) an occupancy estimation method called ALSO [2], 2) occupancy masked autoencoder called Occupancy-MAE [34], 3) multi-modality methods including SLidR [38], PPKT [28] and UniPAD [58]. We use the official implementations to pre-train the backbones with same setting as CLAP.
@@ 4.1 Settings
Implementation Details of CLAP. The feature channels for embeddings $\hat{\mathbf{P}}$, $\hat{\mathbf{I}}$, $\hat{\mathbf{F}}$ and prototypes $\mathbf{K}$ are respectively set to $\hat{d_{\text{P}}}=256$, $\hat{d_{\text{I}}}=80$, $\hat{d_{\text{F}}}=512$ and $\hat{d}_{\text{K}}=128$. Sampling number for point cloud and pixel are $N_{\text{L}}=8192$ and $N_{\text{C}}=1024\times N_{\text{cam}}$. The number of sample points along each ray is $N_{\text{ray}}=96$. Warm-up epochs for Curvature Sampling is $N_{\text{warmup}}=4$. We set the number of learnable prototypes and sinkhorn update iterations to $N_{\text{K}}=512$ and $N_{\text{sink}}=3$. The temperature for swapping prediction loss is $\tau=1.0$. The loss weighting parameters are implemented as $\omega_{\text{r}}=2.0$, $\omega_{\text{proto}}=1.0$, $\omega_{\text{sur}}=0.05$, $\omega_{\text{C}}=0.05$, $\omega_{\text{SwAV}}=1.0$, $\omega_{\text{EM}}=0.1$ and $\omega_{\text{GMM}}=0.1$. We use torch.auto_grad() [37] to implement the derivatives in the curvature estimation. For more details about pre-training and fine-tuning, please refer to Appendix 1.

### v3

@@ 3.2 Neural Field and Differentiable Rendering
Inspired by the success of UniPAD (Yang et al., 2024), CLAP applies a differentiable rendering decoder with neural field to conduct the masked-and-reconstruction pre-training. Different from (Yang et al., 2024), an additional surface signed distance field loss is applied to better optimize the scene geometry. Here we first introduce Neural Field, which is the basis for camera images and point clouds rendering, and then discuss the differentiable rendering process on range and RGB values.
@@
[…]
where $f^{\text{tri}}$ is an built-in module implemented in Pytorch (Paszke et al., 2019). Taking the concatenation of location $\mathbf{p}$ and queried feature $\mathbf{f}_{\text{p}}$ as inputs, we predict the signed distance value $s\in\mathbb{R}$ (Chan & Zhu, 2005; Malladi et al., 1995) and color value $c\in\mathbb{R}^{3}$ (Wang et al., 2021a) at $\mathbf{p}$ with $f^{\text{SDF}}$ and $f^{\text{RGB}}$. $f^{\text{RGB}}$ and $f^{\text{SDF}}$ are parameterized by Multi-layer Perceptron.
@@
Differentiable Rendering. Similar to (Mildenhall et al., 2021; Wang et al., 2021a), we first sample $N_{\text{L}}$ or $N_{\text{C}}$ rays at the LiDAR or camera sensor origin $\mathbf{o}$, each of which is described by its normalized direction $\mathbf{d}$ and $\mathbf{o}$. Next, we sample $N_{\text{ray}}$ points following (Wang et al., 2021a) along each ray. Here each point along the ray can be interpreted by $\mathbf{p}=\mathbf{o}+r\mathbf{d}$, where $r$ is the range from the sensor origin to the point $\mathbf{p}$. Thus the sampled point set can be annotated by $\{\mathbf{p}_{n}=\mathbf{o}+r_{n}\mathbf{d}\}_{n=1}^{N_{\text{ray}}}$ and we predict the estimated signed distance value $s_{n}$ and color value $c_{n}$ for them with $f^{\text{RGB}}$ and $f^{\text{SDF}}$. Following (Wang et al., 2021a), we estimate the occupancy value $\alpha_{n}$ for each sampled point,
@@
Here $\Phi_{h}(x)=(1+e^{-hx})^{-1}$ stands for the sigmoid function paired with a learnable scalar $h$. After that, we predict the accumulated transmittance $t_{n}$ similar to (Wang et al., 2021a)
@@
Based on $t_{n}$, we compute an unbiased and occlusion-aware weight $w_{n}=t_{n}\alpha_{n}$ (Wang et al., 2021a) and integrate all samples along the ray to predict the range $\tilde{r}$ or color $\tilde{c}$ along this ray,
@@
For the observed LiDAR points, it is evident that signed distance value should be 0. The loss function for differentiable rendering is a combination of L-1 loss on surface SDF, range and color predictions.
@@ 3.3 Curvature Sampling
In order to make joint unsupervised representation learning feasible, we have to make $N_{\text{L}}\ll N_{\text{P}}$ and $N_{\text{C}}\ll H\cdot W\cdot N_{\text{cam}}$. Intuitively, uniform sampling with range can be used,
Figure 3: Inspiration of Curvature Sampling.
same as “Memory-friendly Ray Sampling” in (Yang et al., 2024). But due to the relatively small sample number compared to the raw inputs ($\sim\frac{1}{100}$), this sampling method brings little improvement against separate pre-training, which is contradictory to our motivation. Hence, we need to sample more informative part of the scene for $\mathcal{L}_{\text{rend}}$. As shown in Figure 3, we are inspired by the observation that surface with higher curvature (surface of a vehicle) generally contains more information as compared to that with lower curvature (road plane). Therefore, we propose the Curvature Sampling for effective sampling. In order to make it more intuitive, we show camera image here and the curvature estimation is actually computed in 3D with the SDF function. For each point $\mathbf{p}$ in the LiDAR point cloud $\mathbf{P}$, we first estimate surface normal by deriving the signed distance function with respect to $\mathbf{p}$
@@
[…]
@@
[…]
@@ 3.4 Prototype Learning
Expectation-Maximization. In order to guide the learnable prototypes to represent parts of the environment, we propose an Expectation-Maximization (Moon, 1996) training scheme to optimize the prototypes. We first project the LiDAR embeddings $\hat{\mathbf{P}}$ and camera embeddings $\hat{\mathbf{I}}$ separately with two projection heads $f^{\text{proj}}_{\text{P}}$ and $f^{\text{proj}}_{\text{I}}$ to the same dimension as $\mathbf{K}$
[…]
@@
In the Expectation step, we compute the probability $\hat{\mathbf{S}}_{\text{P/I}}$ that each prototype is assigned to each embeddings by applying a softmax operation on $\mathbf{S}_{\text{P/I}}$. Then for Maximization step, we expect to maximize the probability of the assignment between one prototype to one specific part of the scene and this is equal to minimize the entropy of the similarity matrix. Thus, the EM loss is computed as
$\mathcal{L}_{\text{EM}}=-\frac{1}{N_{\text{3D}}N_{\text{K}}}\sum_{n=1}^{N_{\text{3D}}}\sum_{m=1}^{N_{\text{K}}}\{\hat{\mathbf{S}}_{\text{P}}^{n,m}log\hat{\mathbf{S}}_{\text{P}}^{n,m}+\hat{\mathbf{S}}_{\text{I}}^{n,m}log\hat{\mathbf{S}}_{\text{I}}^{n,m}\}.$ (15)
Swapping Prototype Prediction. To further explore interaction between modalities, we detach $\mathbf{S}_{\text{P/I}}$ and apply sinkhorn algorithm (Cuturi, 2013) to approximate them to double stochastic matrix in $N_{\text{sink}}$ iterations. We denote the updated matrix as codes $\mathbf{Q}_{\text{P/I}}\in\mathbb{R}^{N_{\text{3D}}\times N_{\text{K}}}$. The swapping prototype prediction loss is computed with a temperature parameter $\tau$, inspired by (Caron et al., 2020),
@@
Gram Matrix Minimization. When training the randomly initialized prototypes, the network might learn a short cut with all prototypes being the same (Caron et al., 2020), which is called collapse. To avoid this, we estimate similarity between prototypes by the gram matrix $\mathbf{G}=\mathbf{K}\mathbf{K}^{\intercal}$ of prototypes $\mathbf{K}$, the dimension of which is $\mathbf{G}\in\mathbb{R}^{N_{\text{K}}\times N_{\text{K}}}$. Finally we minimize the average of the non-diagonal elements of $\mathbf{G}$ in order to avoid collapse
@@
4 Experiments
Unsupervised 3D representation learning for fusion perception aims to pre-train both LiDAR and camera encoders and initialize downstream models with the pre-trained weights to gain performance improvement in downstream tasks. In this section, we design extensive experiments on the popular autonomous driving dataset NuScenes (Caesar et al., 2020) and Waymo (Sun et al., 2020) to demonstrate the effectiveness of CLAP. To begin with, we describe experiment setups in Section 4.1. Next, we show and analyze main results in Section 4.2. Finally, we provide ablation study and visualizations separately in Section 4.3 and 4.5.
4.1 Settings
Datasets. We use the popular autonomous driving dataset NuScenes (Caesar et al., 2020) and Waymo (Sun et al., 2020) to evaluate the performance of CLAP. NuScenes (Caesar et al., 2020) uses one roof LiDAR and six surrounding cameras to collect data. The LiDAR is a 32-beam Velodyne and collecting frequency is 20Hz. The frequency of camera capturing is 12Hz. (Caesar et al., 2020) conducts the synchronization and provides paired data of LiDAR point cloud and camera images. The whole NuScenes dataset contains 1000 scenes collected in Boston and Singapore. Each scene lasts for around 20 seconds and there are a total of 5.5 hours data. Following convention practice from (Caesar et al., 2020; Team, 2020), we divide the whole dataset into training set with 850 scenes and validation set with 150 scenes. Waymo (Sun et al., 2020) uses one top 64-beam LiDAR, 4 corner LiDARs and 5 surrounding cameras to collect point clouds and camera images. Following the same practice in (Sun et al., 2020; Team, 2020), the collected 1000 scenes are split into training set (798 scenes) and validation set (202 scenes). All pre-trainings are conducted on the training set without labels and we conduct downstream 3D object detection training in few-shot setting via uniform sampling of the training data (NuScenes $5\%$ and Waymo $1\%$).
Downstream 3D Object Detectors. For NuScenes, we select the SOTA 3D object detector called BEVFusion (Liu et al., 2023) and for Waymo, we use CenterPoint (Yin et al., 2021). Both BEVFusion and CenterPoint are implemented in the popular code repository for autonomous driving perception called OpenPCDet (Team, 2020). For evaluation metrics, we use average precisions of various categories (APs), mean average precision (mAP) and NuScenes Detection Score (NDS) (Caesar et al., 2020) for NuScenes and mAP (mean accurate precisions) and mAPH (mean accurate precisions with headings) at different difficulty levels (Level-1 and 2) for Waymo (Sun et al., 2020). We follow a similar setting in (He et al., 2019; et al, 2023) to gradually increase training iterations of the from-scratch model until convergence is observed. Here convergence means further increasing training iterations will not improve the performance. Then the number of training iterations is fixed for fine-tuning pre-trained models. This setting avoids the case that pre-training only accelerates convergence and makes sure that pre-training indeed improve the performance of downstream models, that is improving the sample efficiency of the downstream task.
Baseline Pre-training Method for Fusion Perception. We incorporate three kinds of pre-training baseline methods: 1) an occupancy estimation method called ALSO (Boulch et al., 2023), 2) occupancy masked autoencoder called Occupancy-MAE (Min et al., 2023), 3) multi-modality methods including SLidR (Sautier et al., 2022), PPKT (Liu et al., 2021a) and UniPAD (Yang et al., 2024). We use the official implementations to pre-train the backbones with same setting as CLAP.
Table 1: Results for fine-tuning on 5% of training set in NuScenes. “Init.” means the way to initialize models. We guarantee convergence of from-scratch model and fix the training iteration for all fine-tuning experiments. We provide mAP and NDS as an evaluation of the overall performance of different models and highlight the best mAP and NDS with bold font. We also indicate the performance improvement by green color. “C.V.”, “Mot.”, “Bic.”, “Ped.” and “T.C.” are abbreviations for Construction Vehicle, Motorcycle, Bicycle, Pedestrian and Traffic Cone. Results are in %.
@@ 4.1 Settings
Implementation Details of CLAP. The feature channels for embeddings $\hat{\mathbf{P}}$, $\hat{\mathbf{I}}$, $\hat{\mathbf{F}}$ and prototypes $\mathbf{K}$ are respectively set to $\hat{d_{\text{P}}}=256$, $\hat{d_{\text{I}}}=80$, $\hat{d_{\text{F}}}=512$ and $\hat{d}_{\text{K}}=128$. Sampling number for point cloud and pixel are $N_{\text{L}}=8192$ and $N_{\text{C}}=1024\times N_{\text{cam}}$. The number of sample points along each ray is $N_{\text{ray}}=96$. Warm-up epochs for Curvature Sampling is $N_{\text{warmup}}=4$. We set the number of learnable prototypes and sinkhorn update iterations to $N_{\text{K}}=512$ and $N_{\text{sink}}=3$. The temperature for swapping prediction loss is $\tau=1.0$. The loss weighting parameters are implemented as $\omega_{\text{r}}=2.0$, $\omega_{\text{proto}}=1.0$, $\omega_{\text{sur}}=0.05$, $\omega_{\text{C}}=0.05$, $\omega_{\text{SwAV}}=1.0$, $\omega_{\text{EM}}=0.1$ and $\omega_{\text{GMM}}=0.1$. We use torch.auto_grad() (Paszke et al., 2019) to implement the derivatives in the curvature estimation. More details about pre-training and fine-tuning are provided in Appendix A.